### v1

@@ 4.1 Settings
Datasets. We evaluated EBBS on two popular benchmark datasets for zero-shot machine translation: IWSLT (Cettolo et al., 2017), which contains 4 languages (with English) and 6 zero-shot directions; and Europarl v7 (Koehn, 2005), which contains 9 languages and 56 zero-shot directions.
@@
Model architecture. We replicated Liu et al. (2021) and trained a multilingual system as our base model. Specifically, we had a 5-layer encoder–decoder Transformer for IWSLT, but used an 8-layer Transformer for Europarl to accommodate more training data and languages. We also followed Liu et al. (2021) and removed residual connections in certain middle layers of the encoder (2nd layer for IWSLT; 2nd and 4th layers for Europarl). Liu et al. (2021) show weakening positional correspondence between layers can significantly improve zero-shot performance, which was also observed in our replication experiments.
More details about our experimental setup can be found in Appendix B.
@@ 4.2 Competing Methods
We compare our EBBS against the following methods, including pivoting and different ensembles.
English pivoting. Pivoting is commonly applied for translation (Wu and Wang, 2007, 2009; Vamvas and Sennrich, 2022), and we used English as the pivot in our setting because we have parallel data for translations both from and to English.
Word-level averaging ensemble. Averaging is one of the most widely used ensemble techniques in text generation (Sennrich et al., 2016a; Freitag et al., 2017; Shanbhogue et al., 2023). Essentially, the ensemble components’ probabilities are first averaged before being fed to the standard beam search (Algorithm 2).
@@
Sequence-level voting ensemble. Minimum Bayes risk decoding is originally designed as a single-model decoding algorithm, where it selects a sequence from a set of beam search results based on similarity (Eikema and Aziz, 2020; Müller and Sennrich, 2021). Here, we use it as a sequence-level ensemble technique, where the candidates are the output sequences from different ensemble components. Let $C=\{\mathbf{y}_{1},\cdots,\mathbf{y}_{K}\}$ be the set of candidate outputs given by $K$ ensemble components. The best output is selected as
@@ 4.3 Results and Analysis
IWSLT # Model Average it-nl it-ro nl-it nl-ro ro-it ro-nl
1 Direct translation (Liu et al., ${}^{\dagger}$ 17.7 18.5 17.8 17.9 15.5 19.6 16.8
[…]
@@
4 Word-level averaging ensemble 16.52 16.48 16.49 17.53 13.80 19.07 15.77
[…]
6 Sequence-level voting ensemble (MBR) 16.66 16.54 16.51 17.72 13.64 19.58 15.98
[…]
Europarl # Model Average da-de da-es da-fi da-fr da-it da-nl da-pt
1 Direct translation (Liu et al., ${}^{\dagger}$ 26.9 24.2 33.1 18.1 30.6 26.1 26.3 29.9
[…]
4 Word-level averaging ensemble 27.76 26.21 33.64 18.88 31.09 26.71 27.40 30.37
[…]
@@
Table 1: Main results on the IWSLT and Europarl datasets. The best results are in bold, and the second best results are underlined. ${}^{\dagger}$ indicates cited results; others were obtained by our experimentation.
@@
Main results. We first replicated the base multilingual model Liu et al. (2021). As shown in Rows 1–2, Table 1, the results are generally close, indicating that the replication is successful and ready for our ensemble research. Further, we tried English pivoting (Row 3), a common zero-shot translation method. In our experiments, we find that it does not outperform direct translation, as pivoting methods may suffer from the error accumulation problem since it requires two steps of translation.
@@
We then compare different ensemble techniques, including the proposed EBBS and alternative ensemble methods. Specifically, we notice that IWSLT contains four languages (with English); thus we have two available pivoting directions (excluding source and target), which, along with direct translation, are our three ensemble components. For Europarl, it contains nine languages; for performance and efficiency concerns (to be shown in Figure 2), we also consider three translation paths as our ensemble components: direction translation, English pivoting, and a second pivot. [^2] We use the first available language in the order of Spanish, German, and French. For example, Spanish-to-German translation will have to use French as the pivot. These languages are chosen because they have the most content on the Internet according to the Web Technology Surveys (https://w3techs.com/technologies/overview/content_language), allowing us to better simulate real-world applications.
@@
We start by studying the standard ensemble technique of word-level averaging (Row 4), which has been used in previous translation research (Freitag et al., 2017). In particular, our averaging ensemble performs worse than direct translation on IWSLT, but is slightly better on Europarl. Our results are not consistent to Freitag et al. (2017), who show a word-level averaging ensemble of random seeds can improve performance. This is understandable because models trained with random seeds exhibit similar behavior, and averaging their probabilities achieves a denoising effect. However, our ensemble components are drastically different in terms of their strengths and expertise, because they learn diverse translation patterns from different languages. Thus, word averaging fails to improve translation quality in our setting.
@@
Alternatively, voting ensembles can also be applied, at either the word level or the sequence level. As seen, word-level voting is not effective, as it does not achieve significant improvements (Row 5). This is expected because the voted words (top predictions) by the ensemble components may not overlap due to the large vocabulary size. In such cases, the algorithm defaults to choosing the word with the highest probability, causing the ensemble to follow the most peaked distributions.
@@
Sequence-level voting should also be done in a soft manner, and minimum Bayes risk (MBR) decoding can be thought of as using a Bayes risk to softly “vote” the candidate outputs. As seen from Row 6, such a method works relatively well on Europarl, achieving the second-highest performance across all competing methods; however, it works poorly on the IWSLT dataset. The main drawback of sequence-level voting is that it can only select one of the ensemble components’ outputs. This may not work well when the individual ensemble components are weak, as in our zero-shot setting, and especially with the IWSLT dataset due to a relatively small training set. Such a selective sequence-level ensemble cannot integrate different expertise of its components during generation.
@@
IWSLT Direct 17.28 2 4
Ensemble 18.23 4 2
Europarl Direct 27.85 4 52
Ensemble 28.44 52 4
Overall Direct 26.83 6 56
Ensemble 27.45 56 6
@@
Table 2: Performance statistics on all 66 zero-shot directions in both datasets. $p$-value is given by a two-sided binomial test.
@@
Unlike existing ensemble methods, our EBBS algorithm achieves the highest performance in most directions on both datasets. Notice that Europarl contains 56 zero-shot directions; due to the space limit, we can only present in Table 1 the first seven directions based on the order provided by the dataset. Table 2 further shows a pairwise comparison against direct translation (a strong baseline in our experiment) in all zero-shot directions. As seen, EBBS achieves higher performance in 56 out of 62 cases across two datasets, showing strong statistical evidence for its effectiveness, with a $p$-value of 3e-11 in a two-sided binomial test.
Voting scheme BLEU BLEU1 BLEU2 BLEU3 BLEU4
None (beam search) 25.33 56.32 30.08 19.01 12.78
Total-sum 25.27 56.67 30.27 19.07 12.65
[…]
Top-$Z$ sum (ours) 26.10 57.07 31.00 19.76 13.28
Table 3: Comparison of different ensemble variants. We used Italian-to-Dutch translation in the Europarl dataset as a testbed due to the limit of time and resources.
Analysis of voting methods in EBBS. In our EBBS algorithm, the lower-level beams are synchronized into a shared upper-level beam by voting. Specifically, EBBS uses a sum-voting mechanism, where we add the ensemble components’ probabilities for each appearance of a candidate in the lower-level beam, shown in Eqn. (3). Here, we analyze a few alternative voting methods for EBBS.
If EBBS adopts total-sum voting, it still uses lower-level beams to find candidates, but adds all components’ probabilities together. This is equivalent to applying the common averaging ensemble to the top-$Z$ candidates. However, it differs from our approach, because in total-sum voting, a component will vote even if the candidate does not appear in its own lower-level beam; the probability after voting in Eqn. (3) is substituted with $\frac{1}{K}\sum_{k}p_{k}(\mathbf{y}|\mathbf{x})$. As shown in Table 3, EBBS with total-sum voting performs worse than direct translation, suggesting the importance of ignoring the components whose lower-level beam does not contain the candidate. This is analogous to nucleus sampling Holtzman et al. (2019), where the long tail of a distribution is mainly noise and should be ignored.
Other voting schemes that EBBS may use include 0/1 voting and max voting. The former selects the candidates that appear most in the lower-level beams, disregarding the probability values (unless for ties); the latter chooses the maximum probability across the lower-level beams, which gives preference to sequences through a maximization bias (Hasselt, 2010; van Hasselt et al., 2016). As seen in Table 3, EBBS performs relatively well with both of these voting schemes, achieving a decent improvement over the baseline approach; however, their performance is consistently worse than our sum voting scheme.
Overall, our bi-level beam search ensemble is effective with different voting schemes (except for the total-sum voting), and our sum voting works the best among these variants.
Method BLEU BLEU1 BLEU2 BLEU3 BLEU4
Direct 25.33 56.32 30.08 19.01 12.78
Pivoting (en) 25.08 56.76 30.29 19.06 12.66
Pivoting (es) 24.40 55.38 29.08 18.22 12.09
Pivoting (pt) 24.34 55.46 29.02 18.13 12.02
Pivoting (fr) 24.20 55.41 29.02 18.00 11.84
Pivoting (de) 23.65 55.33 28.69 17.67 11.54
Pivoting (da) 23.12 54.81 27.96 17.12 11.12
Pivoting (fi) 20.74 53.54 26.10 15.43 9.79
[…]
Top-3 ensemble 26.10 57.07 31.00 19.76 13.28
@@
Table 4: The performance of individuals and their ensembles for Italian-to-Dutch translation on Europarl.
[…]
Notably, EBBS can achieve a solid improvement (~0.6 BLEU points) over the baseline despite including weak ensemble components (e.g., Finnish pivoting, which is almost 5 BLEU points lower than direct translation). If we limit the ensemble components to the strong ones (top-3), EBBS can achieve an even higher improvement of ~0.8 BLEU points. This analysis highlights the robustness of our EBBS algorithm.
@@
Figure 2 shows the trends of incremental ensembles. If we add best pivot directions, the performance peaks at three ensemble components; interestingly, the inclusion of weaker components does not affect EBBS much. On the other hand, adding the worst pivot translation leads to an immediate drop of 1.6 BLEU points, which then largely recovers with the second pivot. This is reasonable because the worst pivot (Finnish) is 4.6 BLEU points lower than direct translation, and EBBS cannot decide on which of the two ensemble components to trust; despite this, the performance of EBBS is still much better than the average performance of the components. With a second pivot, it gives a third “opinion” when the first two components “disagree.” The performance continues to rise if more and stronger components are added. In fact, our ensemble even surpasses the baseline with 4 weakest pivot translations, each of which is at least 1 BLEU point lower than the baseline. This demonstrates that EBBS is flexible and works well with both strong and weak ensemble components.
Distillation results. We apply distillation techniques to our ensemble outputs to speed up inference. We follow the same ensemble setting (three components) as in Table 1, but only focus on Italian-to-Dutch translation because generating the training samples for all 66 zero-shot directions (each containing more than 100K samples) requires significant computing resources.
We compare our EBBS-based distillation with two alternatives. A straightforward way to distill from multiple teachers (here, translation paths) is to take the union of teachers’ outputs for training. We also include self-distillation, which does not utilize different translation paths but generates samples by standard beam search for self-training.
As seen in Table 5, self-distillation improves zero-shot translation on both IWSLT and Europarl. This is because the base model has never learned from parallel data between Italian and Dutch, and the self-distillation process enables the multilingual model to further refine its translation capabilities in this zero-shot direction. Similarly, we observe that union distillation also improves performance to some degree, but it is not better than self-distillation. This is consistent with the challenges of one-to-many learning seen in previous work (Wen et al., 2023a, b). In our case, learning from the outputs of multiple teachers in a straightforward manner may confuse the student, failing to consolidate and transfer the teachers’ knowledge.
In comparison, EBBS-based distillation consistently outperforms both self-distillation and union distillation, achieving near-EBBS performance on Europarl. For IWSLT, our model even outperforms EBBS, as IWSLT is a small dataset and the high-quality outputs generated by EBBS can largely improve the base model (no distillation) by 2.7 points.
Method BLEU BLEU1 BLEU2 BLEU3 BLEU4
IWSLT EBBS 19.52 51.87 25.12 13.88 8.02
[…]
Union distillation 17.80 49.21 23.01 12.51 7.10
EBBS distillation 20.13 53.20 26.06 14.33 8.26
Europarl EBBS 26.10 57.07 31.00 19.76 13.28
Direct No distillation 25.33 56.32 30.08 19.01 12.78
Self-distillation 25.44 56.54 30.28 19.13 12.79
[…]
EBBS distillation 25.92 56.76 30.68 19.57 13.24
Table 5: Comparison of various distillation methods for Italian-to-Dutch translation.
We provide an efficiency analysis in Appendix D, showing that the inference time of EBBS is linear to the number of ensemble components. Overall, our EBBS-based distillation achieves a significant speed up while maintaining, or even improving, translation quality.
Additional results. We show additional results in the appendices. C: Analysis of beam size, D: Analysis of inference efficiency, E: Entropy of distilled models, and F: Case study.

### v2

@@ 4.1 Settings
We evaluated EBBS on two popular benchmark datasets for zero-shot machine translation: IWSLT (Cettolo et al. 2017), which contains 4 languages (with English) and 6 zero-shot directions; and Europarl v7 (Koehn 2005), which contains 9 languages and 56 zero-shot directions.
@@
We replicated (Liu et al. 2021) and trained a multilingual translation system as our base model. Specifically, the neural architecture in (Liu et al. 2021) is a 5-layer encoder–decoder Transformer for IWSLT, but has 8 layers for Europarl to accommodate more training data and languages.
For EBBS, we used a beam size of five for both upper- and lower-level beams. In our experiment, we implemented standard beam search for comparison, where we also used a beam size of five, following the common practice (Meister, Cotterell, and Vieira 2020). A comprehensive beam analysis can be found in our appendix.
@@ 4.2 Competing Methods
We comprehensively compare our EBBS with direct/pivot translation and other ensemble methods.
Direct/pivot translation. For direct translation, we applied beam search on the multilingual model to translate in unseen directions. For pivot translation (Wu and Wang 2007, 2009; Vamvas and Sennrich 2022), we used English as the pivot because we have parallel data for translations both from and to English.
Word-level averaging ensemble. Averaging is one of the most widely used ensemble techniques in text generation (Sennrich, Haddow, and Birch 2016a; Freitag, Al-Onaizan, and Sankaran 2017; Shanbhogue et al. 2023). Essentially, the ensemble components’ probabilities are first averaged before being fed to the standard beam search.
@@
Sequence-level voting ensemble. Minimum Bayes risk (MBR) decoding is originally designed as a single-model decoding algorithm, where it selects a sequence from a set of beam search results based on similarity (Eikema and Aziz 2020; Müller and Sennrich 2021). Here, we use it as a sequence-level ensemble technique, where the candidates are the output sequences from different ensemble components. Let $C=\{\mathbf{y}_{1},\cdots,\mathbf{y}_{K}\}$ be the set of candidate outputs given by $K$ ensemble components. The best output is selected as
@@ 4.3 Results and Analysis
IWSLT # Method Average it-nl it-ro nl-it nl-ro ro-it ro-nl
1 Direct translation (Liu et al. 2021) 17.7 18.5 17.8 17.9 15.5 19.6 16.8
2 Direct translation (our replication) 17.29 17.46 17.48 18.23 14.63 19.65 16.26
@@
4 Word-level averaging ensemble 17.28 17.29 17.44 18.33 14.65 19.69 16.30
5 Word-level voting ensemble 16.99 17.58 16.38 17.78 14.13 19.21 16.84
6 Sequence-level voting ensemble (MBR) 16.72 16.64 16.53 17.83 13.74 19.48 16.08
7 EBBS (ours) 18.24 19.52 17.09 19.06 14.58 20.75 18.45
8 Direct w/ EBBS distillation (ours) 18.92 19.86 18.80 19.73 15.39 21.23 18.48
Europarl # Method Average da-de da-es da-fi da-fr da-it da-nl da-pt
1 Direct translation (Liu et al. 2021) 26.9 24.2 33.1 18.1 30.6 26.1 26.3 29.9
2 Direct translation (our replication) 27.74 26.24 33.64 18.95 31.01 26.58 27.36 30.38
3 Pivoting (en) 27.69 25.17 33.87 18.70 31.44 27.12 26.75 30.79
4 Word-level averaging ensemble 27.76 26.13 33.72 18.91 31.01 26.67 27.39 30.50
5 Word-level voting ensemble 27.45 25.76 33.24 18.39 30.96 26.83 26.63 30.37
6 Sequence-level voting ensemble (MBR) 27.90 25.90 33.95 19.15 31.50 27.15 27.09 30.55
@@
Table 1: Main results of BLEU scores on IWSLT and Europarl. The best results are in bold; the second best results are underlined. indicates cited results; others were obtained by our experimentation.
@@
Main results. Our experiment starts by a replication of the base multilingual model (Liu et al. 2021). As shown in Rows 1–2, Table 1, the results are generally close, indicating that our replication is successful and ready for ensemble research. Further, we tried English pivoting (Row 3), a common zero-shot translation method. In our experiments, we find that it does not outperform direct translation, as pivoting methods may suffer from the error accumulation problem due to two-step translation.
@@
We then compare different ensemble techniques, including our proposed EBBS. We notice that IWSLT contains four languages (with English); thus we have two available pivoting directions (excluding source and target), which, along with direct translation, are our three ensemble components. For Europarl, it contains nine languages; for performance and efficiency concerns (to be shown in Figure 2), we also consider three translation paths as our ensemble components: direction translation, English pivoting, and a second pivot. [^3] We use the first available language in the order of Spanish, German, and French. For example, Spanish-to-German translation will have to use French as the pivot. These languages are chosen because they have the most content on the Internet according to the Web Technology Surveys (https://w3techs.com/technologies/overview/content˙language).
@@
We study the common ensemble technique of word-level averaging (Row 4), which has been used in previous translation research (Freitag, Al-Onaizan, and Sankaran 2017). As we can see, the averaging ensemble performs similarly to direct translation on both datasets. Our zero-shot results are different from (Freitag, Al-Onaizan, and Sankaran 2017), which shows a word-level averaging ensemble of random seeds can improve performance in the supervised setting. This is because models trained with different random seeds exhibit similar behavior, and averaging their probabilities achieves a denoising effect. However, our ensemble components differ drastically in terms of their strengths and expertise due to the different translation paths (direct and pivot translations). Thus, word averaging fails to improve translation quality in our setting.
@@
Alternatively, voting ensembles can also be applied, at either the word level or the sequence level. As seen, word-level voting is not effective, as it is worse than direct translation on both datasets (Row 5). This is expected because the voted words (top predictions) by the ensemble components may not overlap due to the large vocabulary size. In such cases, the algorithm defaults to choosing the word with the highest probability, causing the ensemble to follow the most peaked distributions.
@@
Sequence-level voting should also be done in a soft manner, and minimum Bayes risk (MBR) decoding can be thought of as using a Bayes risk to softly “vote” the candidate outputs. As seen from Row 6, such a method works relatively well on Europarl, achieving the second-highest performance across all ensemble methods; however, it works poorly on the IWSLT dataset. The main drawback of sequence-level voting is that it can only select one of the ensemble components’ output. This may not work well when the individual ensemble components are weak, especially with the small IWSLT dataset. Such a selective sequence-level ensemble cannot integrate different expertise of its components during generation.
@@
IWSLT Direct translation 17.29 2 4
EBBS (ours) 18.24 4 2
Europarl Direct translation 27.85 4 52
EBBS (ours) 28.44 52 4
Overall Direct translation 26.83 6 56
EBBS (ours) 27.45 56 6
@@
Table 2: Pairwise comparison on all 62 zero-shot directions in both datasets. The $p$-value is given by a two-sided binomial test.
@@
Unlike existing ensemble methods, our EBBS algorithm achieves higher performance in most directions on both datasets. Noticing that Europarl contains 56 zero-shot directions, we could only present in Table 1 the first seven directions based on the order provided by the dataset, due to the space limit. Table 2 further shows a pairwise comparison against direct translation (a strong baseline in our experiment) in all zero-shot directions. As seen, EBBS achieves higher performance in 56 out of 62 cases across two datasets, showing strong statistical evidence for its effectiveness, with a $p$-value of 3e-11 in a two-sided binomial test.
We also evaluate EBBS-based distillation (Row 8, Table 1). Again, since Europarl has 56 zero-shot directions, we follow the standard practice (Fan et al. 2021) and select a subset of directions, namely, Danish to other languages, to save computational cost. As seen in Row 8, EBBS-based distillation consistently achieves the highest performance in all directions (except for Danish-to-Dutch translation). This shows that an EBBS-distilled model can outperform EBBS, which is not surprising because learning can smooth out the noise of various heuristics (Deshmukh et al. 2021; Jolly et al. 2022), such as the ensemble algorithm in our scenario. Importantly, EBBS-based distillation achieves significantly higher translation quality with no inference overhead compared with direct translation.
Distillation analysis. We compare EBBS-based distillation with other distilling methods. Here, we only focus on Italian-to-Dutch [^4] We could only afford one translation direction for this analysis, because we need to train different models for all competing distilling methods. This differs from Table 1, where we follow previous work and perform EBBS-based distillation for Danish to other languages. We chose Italian-to-Dutch translation here, because it is the first one in IWSLT, conveniently also available in Europarl. translation to save computational cost.
In particular, we consider two alternative distilling methods: direct and union distillation. Direct distillation finetunes the multilingual model with its own predictions based on direct translation. Union distillation, on the other hand, takes the union of the teachers’ outputs (direct and pivot translations) for training, which is under a controlled experimental setup, because it uses exactly the same translation paths as our EBBS-based distillation.
As seen in Table 3, both direct and union distillation marginally improve the performance compared with no distillation. Intriguingly, learning from the union of multiple teachers is not necessarily better than learning from the best teacher (namely, direct translation). This is because multiple teachers may provide conflicting training signals and confuse the student model.
On the contrary, our EBBS-based distillation consistently outperforms direct and union distillation on both datasets. This shows that our ensemble-then-distill approach is able to consolidate the knowledge of multiple teachers to better train the student model.
Dataset Method BLEU BLEU1 BLEU2 BLEU3 BLEU4 chrF2++ TER COMET
IWSLT EBBS 19.52 51.87 25.12 13.88 8.02 45.63 71.36 0.7341
Direct Translation No distillation 17.46 50.49 23.01 12.01 6.66 43.73 72.02 0.7088
Direct distillation 18.10 50.37 23.53 12.63 7.17 44.48 72.86 0.7144
Union distillation 17.80 49.21 23.01 12.51 7.10 44.93 75.92 0.7221
EBBS distillation 20.13 53.20 26.06 14.33 8.26 46.46 69.28 0.7428
Europarl EBBS 26.10 57.07 31.00 19.76 13.28 52.75 65.63 0.8340
Direct Translation No distillation 25.33 56.32 30.08 19.01 12.78 52.32 66.56 0.8276
Direct distillation 25.44 56.54 30.28 19.13 12.79 52.61 66.34 0.8286
Union distillation 25.53 56.58 30.34 19.18 12.91 52.63 66.27 0.8282
EBBS distillation 25.92 56.76 30.68 19.57 13.24 52.73 66.04 0.8307
Table 3: Comparison of various distilling methods for Italian-to-Dutch translation. ↑/↓The higher/lower, the better.
Method BLEU chrF2++ TER COMET
Direct translation 25.33 52.32 66.56 0.8276
Pivoting (en) 25.08 51.92 66.24 0.8322
Pivoting (es) 24.40 51.71 67.91 0.8192
Pivoting (pt) 24.34 51.61 67.68 0.8191
Pivoting (fr) 24.20 51.61 67.84 0.8208
Pivoting (de) 23.65 50.70 67.89 0.8157
Pivoting (da) 23.12 50.36 69.00 0.8156
Pivoting (fi) 20.74 48.11 70.59 0.8051
Our EBBS 26.10 52.75 65.63 0.8340
@@
Table 4: The performance of direct/pivot translation and our EBBS for Italian-to-Dutch translation on Europarl.
Further, the analysis suggests that our EBBS-distilled model achieves a speedup of multiple times compared with EBBS, because after distillation the model is used by direct translation. This is a significant result, because our EBBS-based distillation not only speeds up the EBBS ensemble approach, but also improves the translation quality of EBBS as shown in Row 8, Table 1.
Analysis of ensemble components. In Table 4, we analyze the ensemble components to better understand our ensemble technique for zero-shot machine translation. As seen, direct translation is an effective approach, which is consistent with previous literature (Fan et al. 2021; Liu et al. 2021). English pivoting achieves higher performance for some metrics but lower for others; it is not conclusively better than direct translation, probably because of the error accumulation problem. Pivoting through non-English languages degrades the performance to a large extent because lacking supervision along the pivoting path leads to two steps of zero-shot translation. EBBS, on the other hand, combines the strengths of individual components and consistently outperforms them in all metrics.
@@
Figure 2 shows the trends of incremental ensembles. If we add the best pivot directions, the performance peaks at three ensemble components; interestingly, the inclusion of weaker components does not affect EBBS much. On the other hand, adding the worst pivot translation at the beginning leads to an immediate drop of 1.6 BLEU points, which then largely recovers with the second pivot. This is reasonable because the worst pivot (Finnish) is 4.6 BLEU points lower than direct translation, and EBBS cannot decide on which of the two ensemble components to trust; despite this, the performance of EBBS is still much better than the average performance of the components. With a second pivot, there is a third “opinion” when the first two components “disagree.” The performance continues to rise if more and stronger components are added. In fact, our ensemble even surpasses the baseline with 4 weakest pivot translations, each of which is at least 1 BLEU point lower than the baseline. This demonstrates that EBBS is flexible and works well with both strong and weak ensemble components.
Appendix. We present additional details and results in the appendix:
  •  A.
    Beam search,
  •  B.
    Experimental details,
  •  C.
    Analysis of inference efficiency,
  •  D.
    Average performance across tasks,
  •  E.
    Analysis of beam size,
  •  F.
    Entropy of distilled models,
  •  G.
    Analysis of voting methods in EBBS, and
  •  H.
    Case study.